OpenStudy (anonymous):

How not to end up with [infinity/infinity]? \[\lim_{x \rightarrow -\infty} \frac{ x-4 }{ x }\]

OpenStudy (anonymous):

well......646463688987654357

OpenStudy (luigi0210):

Do you know L'hopital?

OpenStudy (anonymous):

Is L'hopital dividing by the largest variable?

OpenStudy (anonymous):

or something along the lines?

OpenStudy (luigi0210):

Or if you want, you can divide everything by x: \(\Large \frac{x-4}{x}*\frac{1/x}{1/x}\)

OpenStudy (anonymous):

like so? \[\frac{ x/x - 4/x }{ x/x } = \frac{ 1-4/x }{ 1 }\]

OpenStudy (luigi0210):

No, L'hopital's rule is if you get a case like \(\frac{\infty}{\infty}\) or something you are allowed to take the derivative and find the limit again

OpenStudy (anonymous):

and then the limit of 4/x is 0

OpenStudy (luigi0210):

Right, so you just end up with 1 ~

OpenStudy (anonymous):

So the final answer would be 1?

OpenStudy (anonymous):

Ah nice.

OpenStudy (luigi0210):

If you did L'hopital you would of gotten 1 too :P

OpenStudy (luigi0210):

If you ever wanna use/know about L'Hopital's: http://tutorial.math.lamar.edu/Classes/CalcI/LHospitalsRule.aspx

OpenStudy (luigi0210):

Good luck!

OpenStudy (anonymous):

Will check it out, thanks for the help!